OpenStudy (anonymous):

What is the y-intercept of the graph of 3x - 5y = 10? A. 2 B. -2 C. 5 D. 3

OpenStudy (dmndlife24):

y = mx + b b = y-int.

OpenStudy (dmndlife24):

y = 3/5x - 2 so B. is the answer

OpenStudy (anonymous):

B